1 Animals

Animals can eat other animals [REF 0] or plants [REF 1].

2 Plants

Plants are mostly green and are eaten by Animals [REF 0].

3 Books and Writings

Books and writings are general not read by animals [REF 0] or plants [REF 1].

4 Humans

Humans can read books and writings [REF 2].

5 Cross-References

0 Animals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1 Plants . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 2 Books and Writings . . . . . . . . . . . . . . . . . . . . . . . . . 3